# Recurrence Relations and Dispersive Techniques for Precision Multi-Loop Calculations

### v1

@@ Acknowledgements.
$F_{1}$ $=$ $-\frac{\sqrt{-\Delta}}{k^{2}}\left[\ln\left(z_{1}z_{2}\right)+2\rm{i}\pi\right]\;,$ (101)
$F_{2}$ $=$ $-\frac{\sqrt{-\Delta}}{2k^{2}}\left[2{\mbox{Li}}_{2}\left(\frac{1}{z_{1}}\right)+2{\mbox{Li}}_{2}\left(\frac{1}{z_{2}}\right)-\frac{{}_{2}}{{}^{3}}\pi^{2}\!+\!\frac{{}_{1}}{{}^{2}}\ln^{2}{z_{1}}\!+\!\frac{{}_{1}}{{}^{2}}\ln^{2}{z_{2}}+{\rm i}\pi\left(\ln{z_{1}}\!+\!\ln{z_{2}}\right)\right].\hskip 19.91692pt$ (102)
@@
$\ln\left(\frac{k^{2}}{\Delta}\right)F_{1}\Rightarrow\left[\ln\left(\frac{k^{2}}{-\Delta}\right)-{\rm i}\pi\right]F_{1}\;.$ (103)
@@
${\rm Im}\left[{\rm i}^{-1}J^{(2)}(4-2\varepsilon;1,1)\right]$ $=$ $-\frac{\pi^{2-\varepsilon}\Gamma(1+\varepsilon)}{(1-2\varepsilon)}\frac{\pi\sqrt{-\Delta}}{k^{2}}\Biggl\{1+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}$ (104)
$=$ $-\pi^{2-\varepsilon}e^{-\gamma\varepsilon}\;\frac{\pi\sqrt{-\Delta}}{k^{2}}\Biggl\{1+2\varepsilon+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}\;.\hskip 28.45274pt$
@@
${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]$ $=$ $-2\pi^{1-\varepsilon}\Gamma(1+\varepsilon)\;\frac{\pi}{\sqrt{-\Delta}}\Biggl\{1+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}$ (105)
$=$ $-2\pi^{1-\varepsilon}e^{-\gamma\varepsilon}\;\frac{\pi}{\sqrt{-\Delta}}\Biggl\{1+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}\;.\hskip 28.45274pt$
@@
${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]=-2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\frac{\pi}{\sqrt{-\Delta}}\left(\frac{k^{2}}{-\Delta}\right)^{\varepsilon}\;.$ (106)
@@
${\mbox{Li}}_{2}\left(u\right)=-{\mbox{Li}}_{2}\left(\frac{1}{u}\right)+\frac{{}_{1}}{{}^{3}}\pi^{2}-\frac{{}_{1}}{{}^{2}}\ln^{2}{u}-{\rm i}\pi\ln{u}$
(the sign of the imaginary part needs to be fixed, depending on the causal prescription). Using

### v2

@@ Acknowledgements.
$F_{1}$ $=$ $-\frac{\sqrt{-\Delta}}{k^{2}}\left[\ln\left(z_{1}z_{2}\right)-2\rm{i}\pi\right]\;,$ (104)
$F_{2}$ $=$ $-\frac{\sqrt{-\Delta}}{2k^{2}}\left[2{\mbox{Li}}_{2}\left(\frac{1}{z_{1}}\right)+2{\mbox{Li}}_{2}\left(\frac{1}{z_{2}}\right)-\frac{{}_{2}}{{}^{3}}\pi^{2}\!+\!\frac{{}_{1}}{{}^{2}}\ln^{2}{z_{1}}\!+\!\frac{{}_{1}}{{}^{2}}\ln^{2}{z_{2}}-{\rm i}\pi\ln\left(z_{1}z_{2}\right)\right].\hskip 19.91692pt$ (105)
@@
$\ln\left(\frac{k^{2}}{\Delta}\right)F_{1}\Rightarrow\left[\ln\left(\frac{k^{2}}{-\Delta}\right)+{\rm i}\pi\right]F_{1}\;.$ (106)
@@
${\rm Im}\left[{\rm i}^{-1}J^{(2)}(4-2\varepsilon;1,1)\right]$ $=$ $\frac{\pi^{2-\varepsilon}\Gamma(1+\varepsilon)}{(1-2\varepsilon)}\frac{\pi\sqrt{-\Delta}}{k^{2}}\Biggl\{1+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}$ (107)
$=$ $\pi^{2-\varepsilon}e^{-\gamma\varepsilon}\;\frac{\pi\sqrt{-\Delta}}{k^{2}}\Biggl\{1+2\varepsilon+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}\;.\hskip 28.45274pt$
@@
${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]$ $=$ $2\pi^{1-\varepsilon}\Gamma(1+\varepsilon)\;\frac{\pi}{\sqrt{-\Delta}}\Biggl\{1+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}$ (108)
$=$ $2\pi^{1-\varepsilon}e^{-\gamma\varepsilon}\;\frac{\pi}{\sqrt{-\Delta}}\Biggl\{1+\varepsilon\ln\left(\frac{k^{2}}{-\Delta}\right)+{\cal{O}}(\varepsilon^{2})\Biggr\}\;.\hskip 28.45274pt$
@@
${\rm Im}\left[{\rm i}^{-1}J^{(2)}(2-2\varepsilon;1,1)\right]=2\pi^{1-\varepsilon}\frac{\Gamma(1-\varepsilon)}{\Gamma(1-2\varepsilon)}\;\frac{\pi}{\sqrt{-\Delta}}\left(\frac{k^{2}}{-\Delta}\right)^{\varepsilon}\;.$ (109)
@@
${\mbox{Li}}_{2}\left(u\right)=-{\mbox{Li}}_{2}\left(\frac{1}{u}\right)+\frac{{}_{1}}{{}^{3}}\pi^{2}-\frac{{}_{1}}{{}^{2}}\ln^{2}{u}+{\rm i}\pi\ln{u}.$
Here the sign of the imaginary part is fixed according to the causal prescription ($k^{2}\leftrightarrow k^{2}+{\rm i}0$). Using
@@
Analogous results for special cases of $J^{(2)}(2-2\varepsilon;1,1)$ can be obtained by using Eq. (88).